# Ivan Gatica and Alfonso Bustamante

### v1

@@ Abstract
In this paper we present a brief study of the $\sigma$-set-$\sigma$-antiset duality that occurs in $\sigma$-set theory and we also present the development of the integer space $3^{A}=\left\langle 2^{A},2^{A^{-}}\right\rangle$ for the cardinals $|A|=2,3$ together with its algebraic properties. We also develop in this article a presentation of some of the properties of fusion of $\sigma$-sets and finally we present the development and definition of a type of equations of one $\sigma$-set variable.
@@ 1 $\sigma$-Sets and $\sigma$-Antisets
As we said before the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [6] this operation is not associative.
@@ 4 $\sigma$-Sets Equations
$R=\{S_{1},S_{2}\}=\{\{2\},\{1,2\}\}.$
To conclude our study of $\sigma$-sets and $\sigma$-antisets we are going to present a conjecture regarding the possible solutions of a particular type of equations of a $\sigma$-set variable.

### v2

@@ Abstract
In this paper we present a brief study of the $\sigma$-set-$\sigma$-antiset duality that occurs in $\sigma$-set theory and we also present the development of the integer space $3^{A}=\left\langle 2^{A},2^{A^{-}}\right\rangle$ for the cardinals $|A|=2,3$ together with its algebraic properties. In this article, we also develop a presentation of some of the properties of fusion of $\sigma$-sets and finally we present the development and definition of a type of equations of one $\sigma$-set variable.
@@ 1 $\sigma$-Sets and $\sigma$-Antisets
As we said before the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [6, 3, 2] this operation is not associative.
@@ 4 $\sigma$-Sets Equations
$S=\{S_{1},S_{2}\}=\{\{2\},\{1,2\}\}.$
Before we conclude our study of $\sigma$-sets and $\sigma$-antisets we are going to present a conjecture regarding the possible solutions of a particular type of equations of a $\sigma$-set variable.
@@
Let us look into the following case, where $M=\{1^{*}\}$ and $N=\{1\}$, so for the equation
$X\oplus\{1^{*}\}=\{1\},$
there is no element $x$ such that $\{x,1^{*}\}=\{1\}$. This lead us into a more general case where $N\subseteq M^{-}$ or $N\supset M^{-}$, being the conjecture false, which take us into establishing the following condition in order for this type of equation being solvable.
Definition 4.10.
A $\sigma$-set equation $X\oplus M=N$ is said to be fusionable if $M\widehat{\cap}N=\emptyset$.
With this, let us conclude with a bounded theorem to find some solutions of the $\sigma$-set equation.
Theorem 4.11.
Let $A$ be an integer $\sigma$-set, $X$ a $\sigma$-set variable of $3^{A}$, and $M,N\in 3^{A}$, then two possible solutions $S=\{S_{1},S_{2}\}$ of the fusionable equation
$X\oplus M=N,$
are $S_{1}=N\oplus R^{-}$ and $S_{2}=R^{-}$, where $R:=M\oplus N^{-}.$
Proof.
For the first solution $S_{1}$ we have that
$S_{1}$ $=$ $N\oplus R^{-}$
$=$ $N\oplus(M\oplus N^{-})^{-}$
$=$ $N\oplus(N\oplus M^{-})$
$=$ $N\oplus M^{-},$
where $S_{2}=(M\oplus N^{-})^{-}=N\oplus M^{-}$ because of the result iteration seen above. Hence both results are actually a fusion solution for $X\oplus M=N,$ where $S_{2}=R^{-}$ is an exact solution and $S_{1}=N\oplus R^{-}$ is an intersected rest solution. Because of $M\widehat{\cap}N=\emptyset$ (Definition 4.10) as the equation $X\oplus M=N$ is fusionable, both $S_{1}\oplus M$ and $S_{2}\oplus M$ will be fusionable into another $\sigma$-set $N$. ∎
As we looked above, the solution space is reduced such that the solutions are indeed $N\oplus M^{-},$ being by consequence possible solutions for the fusionable equation $X\oplus M=N.$
Example 4.12.
Let $A=\{1,2,3,4,5,6\}$ be an integer $\sigma$-set, $M=\{1,2,3^{*},4^{*},5,6^{*}\}$ and $N=\{1,2\}$, then the equation of a $\sigma$-set variable
$X\oplus\{1,2,3^{*},4^{*},5,6^{*}\}=\{1,2\},$
which is fusionable because $M\widehat{\cap}N=\{1,2,3^{*},4^{*},5,6^{*}\}\widehat{\cap}\{1,2\}=\emptyset.$
Now, by using Theorem 4.11, let us first obtain
$R$ $=$ $(M\oplus N^{-})^{-}$
$=$ ($(\{1,2,3^{*},4^{*},5,6^{*}\}\oplus\{1,2\}^{-})^{-}$
$=$ ($(\{1,2,3^{*},4^{*},5,6^{*}\}\oplus\{1^{*},2^{*}\})^{-}$
$=$ ($(\{3^{*},4^{*},5,6^{*}\})^{-}$
$=$ $\{3,4,5^{*},6\},$
so we get $S_{1}=N\oplus R^{-}=\{1,2,3,4,5^{*},6\}$ and $S_{2}=R^{-}=\{3,4,5^{*},6\}$, which can be easily proved that both solutions gives $S_{1}\oplus M=S_{2}\oplus M=N$ as a resulting $\sigma$-set. Hence $S=\{\{1,2,3,4,5^{*},6\},\{3,4,5^{*},6\}\}$ is a solution set for the fusionable equation $X\oplus M=N$.